Joe left a campsite 4 hours after Liz left and they traveled in opposite directions. Liz was in a car that traveled at 48 km/hr and Joe was in a car that traveled at 72 km/hr. How far has each person traveled when are they 1008 km apart?
Answer by mananth(16946)

Joe left a campsite 4 hours after Liz left and they traveled in opposite directions. Liz was in a car that traveled at 48 km/hr and Joe was in a car that traveled at 72 km/hr. How far has each person traveled when are they 1008 km apart?
Liz was in a car that traveled at 48 km/hr
in 4 hours she had traveled 4*48 = 192 km when Joe starts
Distance between them when Joe starts = 1008-192= 816 km
and they traveled in opposite directions.
Add up speeds 72+48=120 km/h

Time Joe takes to be 1008 km away = 816/120= 6.18 hours
Liz travels 48*6,8+192=518.4 km
Joe travels 72*6,8=489,6 km.